OpenStudy (anonymous):

How many moles of N are in .227 grams of N2O?

OpenStudy (paxpolaris):

what's the molar mass of N2O?

OpenStudy (paxpolaris):

14*2 + 16*1 = 44

OpenStudy (paxpolaris):

.227 g of N2O has: .227/44 moles of N2O

OpenStudy (paxpolaris):

since there are 2 N in ever N2O multiply it by 2. to get your answer

OpenStudy (paxpolaris):

\[.227\ g\ N_2O \times {1\ mole\ N_2O \over 44.0\ g\ N_2O}\times {2\ mole\ N \over 1\ mole\ N_2O}\]

OpenStudy (anonymous):

Thank you!